OpenStudy (annie96):

Jane is traveling to Europe this summer with the French club. She plans to bring $300 to spend while she is there. If $90 in U.S. currency is equivalent to 63 euros, how many euros will she receive when she exchanges her money?

OpenStudy (kash_thesmartguy):

\[\frac{ 63 }{ 90 }=\frac{ x }{ 300 }\]\[300*0.7=210\] Euros 210 is your answer. You get it? @Annie96